OpenStudy (nani200):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 2, -4, and 1 + 3i

OpenStudy (tkhunny):

Can you write a polynomial function of minimum degree and Real coefficients that has 2 as its ONLY zero?

OpenStudy (nani200):

im not understanding what youre asking

satellite73 (satellite73):

if \[1+3i\] is a zero, there must be another one do you know it?

satellite73 (satellite73):

the question is not that hard, if you know what to do

OpenStudy (tkhunny):

If there is a zero of "2", there is a factor of (x-2). If there is a zero of "-2", there is a factor of (x-(-4)) = (x+4). That's all that part is. Then there is what satellite asked.

OpenStudy (nani200):

so is the other one 1-3i

OpenStudy (mathmale):

Yes, and so you have at minimum four roots: 2, -4, 1+3i and 1-3i. How would you turn these roots into factors of the desired polynomial?

OpenStudy (nani200):

it would be x-2 and x+4 I think?

satellite73 (satellite73):

those are two factors, yes

satellite73 (satellite73):

ow your harder (but that that hard) job is to find the one with zeros \(1\pm 3i\) do you know how to do that? there is a hard way, an easy way, and a real real easy way you pick

OpenStudy (nani200):

i dont know how to do that

satellite73 (satellite73):

ok one way is to work backwards

satellite73 (satellite73):

put \[x=1+3i\] like you just solved a quadartic equation by completing the square the previous step would be \[x-1=3i\]

satellite73 (satellite73):

squaring both sides gives \[(x-1)^2=(3i)^2\\ (x-1)^2=-9\]

satellite73 (satellite73):

expand to get \[x^2-2x+1=-9\] add 9 to get your quadratic \[x^2-2x+10=0\]

OpenStudy (nani200):

that's the answer?

satellite73 (satellite73):

the other way is to memorize that if one root is \(a+bi\) then the quadratic is \[x^2-2ax+(a^2+b^2)=0\]

satellite73 (satellite73):

that is not your final answer

satellite73 (satellite73):

your final answer is \[(x+4)(x-2)(x^2-2x+10)\]

satellite73 (satellite73):

or whatever you get when you multiply that mess out

OpenStudy (nani200):

ok so i just need to distribute all of that and i should get the answer?

OpenStudy (nani200):

@satellite73 can you help me multiply it out? i got -x^3 + 5x^2 + 2x + 32 but that isnt an answer choice

OpenStudy (nani200):

answer choices are A. x^4 - 2x^2 +36x - 80 B. x^4 - 3x^3 + 6x^2 - 18x + 80 C. x^4 - 9x^2 +36x - 80 D.x^4 - 3x^3 - 6x^2 + 18x - 80

satellite73 (satellite73):

where did the negative leading coefficient come from?

satellite73 (satellite73):

\[(x+4)(x-2)(x^2-2x+10)\] is what you need right? multiplying out is donkey work, let the computer do it

OpenStudy (nani200):

yes thats it

satellite73 (satellite73):

or else you have to first multiply \[(x+4)(x-2)\] and see what you get what do you get?

OpenStudy (nani200):

x^2 + 2x - 8

satellite73 (satellite73):

looks good

satellite73 (satellite73):

now you have to multiply \[(x^2+2x-8)(x^2-2x+10)\]

satellite73 (satellite73):

which is a pain, you have \(3\times 3=9\) multiplications to do, then you have to combine like terms there is a much easier way

OpenStudy (nani200):

i got x^4 + 16 x

satellite73 (satellite73):

not even close

OpenStudy (nani200):

whats the easier way

satellite73 (satellite73):

scroll down to "alternate form" http://www.wolframalpha.com/input/?i=(x%2B4)(x-2)(x%5E2-2x%2B10)

OpenStudy (nani200):

thank you!!!